Find the length of spiral curve T() = ----- 0 < > < 2”
Answer #1

e = TE = de 2 LE --16 --- + 20 20 e LE + 2 2 0 L =j vazo de = fede L=[eo] L= 0 0 L = e- L=e - 1 L = 22.1407